# Assessing the Role of Volumetric Brain Information in Multiple Sclerosis Progression

### v1

@@ Abstract
Keywords: multiple sclerosis; deformation-based morphometry; survival analysis; feature selection; random forest; convolutional neural network
Abbreviations: MS, multiple sclerosis; MRI, magnetic resonance imaging; DBM, deformation-based morphometry; LOCO-MP, leave-one-covariate-out minipatch; RSF, random survival forest; CNN, convolutional neural network; CCDP24, confirmed composite disability progression sustained for at least 24 weeks; T25FW, timed 25-foot walking speed test; 9HPT, 9-hole peg test; EDSS, Expanded Disability Status Scale; S25FW, disability progression in 25-foot walking test sustained for at least 24 weeks; BT25FW, baseline timed 25-foot walking speed test
@@ 1 Introduction
Multiple sclerosis (MS) is a chronic autoimmune disease that affects the central nervous system (CNS) in young adults (Compston and Coles,, 2008). Magnetic resonance imaging (MRI) of the brain is a key tool for diagnosing and monitoring MS progression. Different types of MRI data possess overlapping yet complementary properties and applications. For instance, brain volume measurements derived from T1-weighted MRI provide insights into brain atrophy, which is associated with disability and disease progression (Chard et al.,, 2002; Bermel and Bakshi,, 2006; Lomer et al.,, 2024; Fisniku et al.,, 2008). Deformation-based morphometry (DBM), an image registration-based method (Ashburner et al.,, 1998; Chung et al.,, 2001), provides voxel-level measurements of volume differences across the entire brain which offers finer-grained insights into spatially distributed neurodegeneration (Azevedo et al.,, 2018; Rocca et al.,, 2021). MRI data has proven instrumental in predicting disease progression and plays a critical role in personalizing patient treatment decisions (Rocca et al.,, 2019). Additionally, identifying specific brain regions with volume changes linked to disease progression enhances patient monitoring and facilitates the evaluation of treatment efficacy (Song et al.,, 2022).
@@
Despite recent progress in using T1-weighted MRI data to predict MS progression (Moazami et al.,, 2021; Vázquez-Marrufo et al.,, 2022; Lomer et al.,, 2024; Prathapan et al.,, 2024), several open questions remain regarding how data and methodological choices may affect model reliability. For example, recent work leveraging random survival forests achieved high accuracy in predicting clinical progression over a 10-year period based on changes in T1-weighted MRI-derived features within the first 2 years (Pisani et al.,, 2021). However, baseline T1-weighted MRI data alone demonstrated relatively limited predictive power for 2-year progression (Pellegrini et al.,, 2020). Similarly, it was shown that, while voxel clusters derived from independent components analysis (ICA) applied to 2-year volume change data were predictive of 10-year progression, the voxel cluster values at baseline were similar between progressors and non-progressors (Bergsland et al.,, 2018). Some studies have also noted unclear predictive value of T1-weighted MRI, which can vary between studies and can depend on outcome definitions (Pellegrini et al.,, 2020; Tousignant et al.,, 2019). These issues also persist in deep learning frameworks (Zhang et al.,, 2023; Coll et al.,, 2024). It also remains unclear whether T1-weighted MRI data provides complementary information for progression prediction beyond clinician-measured variables of disease state, some of which are used to define common progression endpoints themselves (Sormani et al.,, 2010; Cadavid et al.,, 2017). Ultimately, there is a lack of literature on detecting baseline volume measurements from T1-weighted brain MRI that are predictive of clinical progression within two years. For broader reviews of data-driven approaches to clinical modeling of MS, we refer readers to Moazami et al., (2021); Vázquez-Marrufo et al., (2022); Lomer et al., (2024); Prathapan et al., (2024).
@@
In this study, we utilize deformation-based morphometry derived from T1-weighted MRI of a Phase III MS clinical trial (Hauser et al.,, 2017; Song et al.,, 2022) to investigate the impact of various methodological choices, particularly those used to identify predictive DBM features, on the reliability of conclusions drawn from this dataset. We first employ a state-of-the-art, model-agnostic feature importance algorithm, adapted using leave-one-covariate-out methods (Lei and Wasserman,, 2014) and minipatch ensembles (LOCO-MP) (Gan et al.,, 2023), to identify brain regions from low-signal T1-weighted MRI baseline data that are predictive of MS progression. While general feature importance metrics can be applied, they often yield inconsistent results when the data are highly correlated and high-dimensional (Agarwal et al.,, 2023). LOCO-MP leverages learning from small feature subsets to isolate the effects of highly correlated regions. When applied to region-based summary statistics of volumetric T1-weighted MRI data, this method identifies regions implicated in MS progression, outperforming standard feature importance approaches.
@@
Next, we apply traditional survival models to examine how different outcome/endpoint definitions, in conjunction with feature selection by LOCO-MP, influence prediction power. Although our progression endpoint (CCDP24, discussed in Section 2.1) is highly censored within the two-year study period, it represents a well-defined and clinically meaningful composite marker of MS progression. To validate the DBM features identified using LOCO-MP, we assess their generalizability across alternative progression endpoints (S25FW, discussed in Section 2.1). This also allows us to compare a single, objective measure of physical disability with the composite measure, the latter involving some degree of clinician discretion (Sormani et al.,, 2010). By using DBM at baseline to identify patient volumetric abnormalities and by applying LOCO-MP to enforce sparsity in the DBM feature set, we achieve improved prediction performance compared to models that use whole-brain T1-weighted MRI data as input. Interestingly, models using selected MRI features alone perform on par with those based solely on conventional (non-DBM) measurements. Combined models incorporating both selected MRI features and conventional measures do not necessarily outperform individual models (those using either selected MRI features or conventional measures alone). Notably, removing baseline conventional variables that define certain endpoints significantly reduces the predictive power of the conventional feature models, underscoring the strong predictive value of the selected MRI features.
@@
We further analyze whether 3D convolutional neural networks (CNNs) using only the LOCO-MP regions perform better than full brain architectures. We develop a 3D CNN architecture, Region CNN, that accepts a set of atlas region tensors as input. While full brain MRI deep transformer models (Dai et al.,, 2021; Yu et al.,, 2023) are unstable and show low prediction accuracy, Region CNN using LOCO-MP identified features performs substantially better. This highlights the importance in filtering out unnecessary regions via LOCO-MP for extracting prognostic signal from CNNs for this cohort.
@@ 2.1 Data and image preprocessing
The data used in this study are from a comparator arm of a phase 3 clinical trial of relapsing multiple sclerosis (OPERA I: NCT01247324), in which the patients were treated with interferon (IFN) $\beta$-1a (44 $\mu$g) three times per week throughout the 96-week treatment period. Our sample consisted of 350 patients. Details on patient selection, MRI acquisition, and clinical assessments are provided in the original report (Hauser et al.,, 2017). Briefly, patients were recruited with an age range of 18 to 55 years; McDonald criteria diagnosis of multiple sclerosis; Expanded Disability Status Scale (EDSS) score between 0 to 5.5; more than 2 documented clinical relapses within the previous 2 years or one clinical relapse within the year before screening; brain MRI evidence of multiple-sclerosis-related abnormalities. Conventional T1-weighted 3D spoiled gradient-recalled echo brain MRI was acquired at baseline, Weeks 24, 48 and 96 (repetition time = 28–30 ms, echo time = 5–11 ms, flip angle = 27–30 deg, 60 oblique axial slices of 1 mm in-plane resolution and 3 mm slice thickness). The primary clinical outcome was defined as a composite measure of disability progression which included three clinical assessments: EDSS, Timed 25-Foot Walk (T25FW, a measure of short distance walking speed), and 9-Hole Peg Test (9HPT, a measure of upper limb function) (Elliott et al.,, 2019). This composite measure was developed to capture a broader aspect of disability in patients with multiple sclerosis. The 24-week composite confirmed disability progression (CCDP24) was defined as progression on any one of the three components (EDSS, T25FW, or 9HPT): an increase of EDSS score from the baseline at least 1.0 point (or 0.5 points if the baseline EDSS score was larger than 5.5) or a 20% minimum threshold change for T25FW and 9HPT. We also consider a 24-week progression outcome based solely on the 25-foot walking score (S25FW) in Section 3.2.2. Table 1 summarizes the data in terms of dimension and number of relevant features.
@@
Since MS progression is influenced by many factors, our models consider a variety of baseline demographic and clinical measurements as well as standard imaging metrics. Broadly speaking, we refer to these features as “conventional features” to distinguish them from DBM-based features. The conventional features serve as a comparison against the DBM features to assess which feature set performs more optimally. We record nine conventional features which include the patient’s age, birth sex, years of onset, weight (in kg), total brain volume, total T2 lesion volume, baseline 25-foot walking test score, baseline 9-hole peg test score, and baseline EDSS, which are generally thought to be related to MS disease progression (Bar-Or et al.,, 2023).
Sample size # conventional features # DBM regions Summary features Outcomes
350 9 51 Median, St. Dev CCDP24, S25FW
Table 1: OPERA I data summary.
@@
Statistical features of regional brain volume were extracted from T1-weighted brain MRI, using a deformation-based morphometry (DBM) pipeline based on the diffeomorphic image registration of advanced normalization tools (Tustison et al.,, 2019).
@@
The Mindboggle atlas was used to identify individual brain regions (Klein et al.,, 2017). In addition, prior probability images of six major brain regions based on the Mindboggle atlas, including CSF, cortical grey matter, deep grey matter, cortical white matter, the brainstem, and the cerebellum, were binarized at a probability threshold of 0.5. A population-specific group template of the T1-weighted brain MRI was constructed by advanced normalization tools with T1-weighted brain MRI images of 171 healthy adults (aged 20–59 years). Before being fed into the pipeline, each individual brain image was preprocessed with the following steps: 1) resampling to an isotropic resolution of 1 mm; 2) N4 bias correction (Tustison et al.,, 2010); and 3) denoising with a non-local algorithm (Manjón et al.,, 2010). Regional labels were mapped from the group template to individual images so that volume change of any predefined region could be measured.
The tabular DBM feature set consists of the median and standard deviation voxel values for each of the 51 atlas regions, resulting in 102 total features. Prior to performing feature selection and modeling, we filter out features with a variance below 0.01 across all patients, reducing the number of DBM features from 102 to 56. No other conventional features were excluded in this filtering step.
@@ 2.2 Feature importance using minipatch ensembles
In high dimensional data sets, variable importance metrics are commonly used to filter out variables that are inconsequential to prediction and retain those with the most predictive value. For random forest models, feature importance is typically summarized by averaging the “out-of-bag” prediction error across trees when that feature is permuted (Ishwaran and Kogalur,, 2019; Breiman et al.,, 2018). However, these metrics may produce misleading results, especially when correlated features are present (Agarwal et al.,, 2023). As such, to identify the most important DBM features and brain regions associated with MS progression, we extend a feature importance method called leave-one-covariate-out minipatch (LOCO-MP) prediction (Lei and Wasserman,, 2014; Yao and Allen,, 2021; Gan and Allen,, 2022; Gan et al.,, 2023) to survival analysis. Our outcome of interest, patient progression, is recorded as a time-to-event pair $(T_{i},C_{i})$, where $T_{i}$ is the event time and $C_{i}\in\left\{0,1\right\}$ is an indicator for whether the patient experiences progression ($C_{i}=1$) or is censored for whatever reason ($C_{i}=0$). LOCO-MP determines feature importance for a certain target feature by aggregating predictions across multiple data subsamples (called “minipatches”) and comparing the average prediction performance across minipatches with and without that feature. Dependencies between features are addressed through aggregating the minipatches, which ensures that each feature’s predictive value is not overshadowed by other correlated features. The procedure is summarized graphically in Figure 2 and outlined below for a target feature $j$:
@@
    Repeat steps 1 and 2 over multiple minipatches.
@@
Because our outcomes of interest are recorded in a time-to-event format, we extend LOCO-MP to survival analysis by utilizing discrete hazard loss functions to assess individual-level error, as discussed further below.
@@
Details of LOCO-MP for survival analysis
Following the notation of Gan et al., (2023), we denote the $N\times M$ data matrix as $X$ and the outcome as $Y$. Let $(X_{i},Y_{i})$ denote the feature-outcome pair for individual $i$, where $X_{i}$ is a $M$-dimensional vector and $Y_{i}=(T_{i},C_{i})$. For every minipatch $k=1\dots K$, we sample $n$ observations $I_{k}$ and $m$ features $F_{k}$ from $X$. Denote this minipatch as $X_{I_{k},F_{k}}$. A prediction model $\hat{\mu}_{k}$ is then trained on $X_{I_{k},F_{k}}$ and predictions are generated for the remaining observations $X_{{-I_{k},F_{k}}}$, which includes features $F_{k}$ and observations outside $I_{k}$. Repeating this over all $K$ minipatches, we compute the individual feature occlusion score for feature $j$ and observation $i$:
where
$\bar{\Delta}_{j}=\frac{1}{N}\sum_{i=1}^{N}\hat{\Delta}_{ij}.$ (1)
The prediction error for survival analysis models is commonly measured as the reciprocal of the overall concordance index (C-index) of risk scores between pairs of observations. However, this metric is not an aggregation of observation-level errors, preventing its direct usage in the LOCO-MP framework. To address this, we utilize discrete hazard loss functions used in deep learning survival analysis models (Zadeh and Schmid,, 2020; Vale-Silva and Rohr,, 2021; Shao et al.,, 2023). Specifically, we evenly divide the survival time scale of the study into $d$ intervals: $[t_{0},t_{1}),\ldots,[t_{d-1},t_{d})$, where $t_{d}$ is the end of the study period. The patient’s event time is now denoted $T_{i}=q\;\text{if and only if}\;T_{i}\in[t_{q},t_{q+1})$.
$h(q\mid X_{i})=\mathbb{P}(T_{i}=q\mid T_{i}\geq q,X_{i}).$
$\ell_{U}(T_{i}=q)=h(q\mid X_{i})~{}S(q-1\mid X_{i}).$
$\ell_{C}(T_{i}>q)=S(q\mid X_{i}).$
Here, $\hat{\mu}_{-i}^{-j}$ encapsulates the collection of conditional hazards that are computed with the underlying minipatch model. Assuming that the patient conditional hazard probabilities are not exactly 0 or 1 over the study period, this loss framework satisfies the necessary assumptions for the theoretical guarantees of LOCO-MP (Gan et al.,, 2023).
LOCO-MP is advantageous over other feature selection techniques due to its ability to provide asymptotic inferential guarantees (Gan et al.,, 2023). Moreover, LOCO-MP can also account for dependencies across features, a common issue in high-dimensional settings like ours – Gan et al., (2023) discuss that, by generating randomly subsampled features across minipatches, LOCO-MP ensures that the predictive value of each feature is not being diminished by other strongly correlated features, since groups of strongly correlated features will not always appear in the same minipatch. The full LOCO-MP procedure and its theoretical guarantees are discussed in Gan et al., (2023).

### v2

@@ Abstract
keywords:
multiple sclerosis; deformation-based morphometry; survival analysis; feature selection; random forest; convolutional neural network
journal: Computational and Structural Biotechnology Journal
[…]
Abbreviations: MS, multiple sclerosis; MRI, magnetic resonance imaging; DBM, deformation-based morphometry; VBM, voxel-based morphometry; LOCO-MP, leave-one-covariate-out minipatch; RSF, random survival forest; CNN, convolutional neural network; CCDP24, confirmed composite disability progression sustained for at least 24 weeks; T25FW, timed 25-foot walking speed test; 9HPT, 9-hole peg test; EDSS, Expanded Disability Status Scale; S25FW, sustained disability progression in 25-foot walking test sustained for at least 24 weeks; BT25FW, baseline timed 25-foot walking speed test
@@ 1 Introduction
Multiple sclerosis (MS) is a chronic autoimmune disease that affects the central nervous system (CNS) in young adults [1]. Magnetic resonance imaging (MRI) of the brain is a key tool for diagnosing and monitoring MS progression. Different types of MRI data possess overlapping yet complementary properties and applications. For instance, brain volume measurements derived from T1-weighted MRI provide insights into brain atrophy, which is associated with disability and disease progression [2, 3, 4, 5]. Identifying specific brain regions with atrophy and volume changes that are linked to disease progression enhances patient monitoring and facilitates the evaluation of treatment efficacy [6]. Early efforts to assess volume changes in MS consisted of using voxel-based morphometry (VBM), a method designated for volume assessment of grey matter regions [7]. VBM demonstrated limited associations with clinical outcomes in MS studies [8] and a majority of VBM studies were cross-sectional and thus unable to track disease-related changes at the individual level. On the other hand, deformation-based morphometry (DBM) [9, 10] provides voxel-level measurements of volume assessment across the entire brain, including grey matter, white matter, and the ventricles. We note that enlargement of the fluid-filled ventricles in the brain typically reflect either local or global brain tissue loss, which cannot be captured by VBM due to its focus on grey matter regions. While VBM often suffers from tissue segmentation errors [10, 11], DBM only relies on image registration, making it a more robust tool to assess voxel-based volume differences.
@@
MRI data has proven instrumental in predicting disease progression and plays a critical role in personalizing patient treatment decisions [12, 13]. Despite recent progress in using T1-weighted MRI data to predict MS progression [14, 15, 4, 16], several open questions remain regarding how data and methodological choices may affect model reliability. For example, recent work leveraging random survival forests achieved high accuracy in predicting clinical progression over a 10-year period based on changes in T1-weighted MRI-derived features within the first 2 years [17]. However, baseline T1-weighted MRI data alone demonstrated relatively limited predictive power for 2-year progression [18]. Similarly, it was shown that, while voxel clusters derived from independent components analysis (ICA) applied to 2-year volume change data were predictive of 10-year progression, the voxel cluster values at baseline were similar between progressors and non-progressors [19]. Some studies have also noted unclear predictive value of T1-weighted MRI, which can vary between studies and can depend on outcome definitions [18, 20]. These issues also persist in deep learning frameworks [21, 22]. It also remains unclear whether T1-weighted MRI data provides complementary information for progression prediction beyond clinician-measured variables of disease state, some of which are used to define common progression endpoints themselves [23, 24]. Ultimately, there is a lack of literature on detecting baseline volume measurements from T1-weighted brain MRI that are predictive of clinical progression within two years. For broader reviews of data-driven approaches to clinical modeling of MS, we refer readers to [14, 15, 4, 16, 25].
@@
In this study, we utilize baseline deformation-based morphometry derived from T1-weighted MRI of a Phase III MS clinical trial [26, 6] to investigate the impact of various methodological choices, particularly those used to identify predictive DBM features, on the reliability of conclusions drawn from this dataset. We first employ a state-of-the-art, model-agnostic feature importance algorithm, adapted using leave-one-covariate-out methods [27] and minipatch ensembles (LOCO-MP) [28], to identify brain regions from low-signal T1-weighted MRI baseline data that are predictive of MS progression. Broadly speaking, LOCO-MP computes feature importance by aggregating predictions from models trained on multiple small, randomly selected subsets of the data (minipatches). We observe that LOCO-MP combined with a survival random forest model identifies features in a more consistent and stable manner than survival random forests alone. This is because LOCO-MP leverages learning from small feature subsets to isolate the effects of highly correlated regions while other methods often yield inconsistent results when the data are highly correlated and high-dimensional [29].
@@
Next, we apply traditional survival models to examine how different outcome/endpoint definitions, in conjunction with feature selection by LOCO-MP, influence prediction power. Although our progression endpoint (CCDP24, discussed in Section 2.1) is highly censored within the two-year study period, it represents a well-defined and clinically meaningful composite marker of MS progression. To validate the DBM features identified using LOCO-MP, we assess their generalizability across alternative progression endpoints (S25FW, discussed in Section 2.1) to compare a single, objective measure of physical disability with the composite measure, the latter involving some degree of clinician discretion [23]. We also verify that our selected features generalize well across a separate patient cohort (see D).
By using DBM at baseline to identify patient volumetric abnormalities and by applying LOCO-MP to enforce sparsity in the DBM feature set, we achieve improved prediction performance compared to models that use whole-brain T1-weighted MRI data as input. Additionally, we compare the model performance of DBM features against models using conventional, non-DBM clinical features (such as age and sex), also measured at baseline. We introduce these conventional features in Section 2.1. Interestingly, models using selected MRI features alone perform on par with those using only the baseline conventional measurements. Combined models incorporating both selected MRI features and conventional measures do not necessarily outperform individual models (those using either selected MRI features or conventional measures alone). Notably, removing the conventional features that define certain endpoints significantly reduces the predictive power of the conventional feature models, underscoring the strong predictive value of the selected MRI features.
@@
We further analyze whether 3D convolutional neural networks (CNNs) using only the LOCO-MP regions perform better than full brain architectures. We develop a 3D CNN architecture, Region CNN, that accepts a set of atlas region tensors as input. While full brain MRI deep transformer models [30, 31] are unstable and show low prediction accuracy, Region CNN using LOCO-MP identified features performs substantially better. This highlights the importance in filtering out unnecessary regions via LOCO-MP for extracting prognostic signal from CNNs for this cohort.
@@ 2.1 Data and image preprocessing
The data used in this study are from a comparator arm of a phase 3 clinical trial of relapsing multiple sclerosis (OPERA I: NCT01247324), in which the patients were treated with interferon (IFN) $\beta$-1a (44 $\mu$g) three times per week throughout the 96-week treatment period. Our sample consisted of 350 patients. Details on patient selection, MRI acquisition, and clinical assessments are provided in the original report [26]. Briefly, patients were recruited with an age range of 18 to 55 years; McDonald criteria diagnosis of multiple sclerosis; Expanded Disability Status Scale (EDSS) score between 0 to 5.5; more than 2 documented clinical relapses within the previous 2 years or one clinical relapse within the year before screening; brain MRI evidence of multiple-sclerosis-related abnormalities. Conventional T1-weighted 3D spoiled gradient-recalled echo brain MRI was acquired at baseline, Weeks 24, 48 and 96 (repetition time = 28–30 ms, echo time = 5–11 ms, flip angle = 27–30 deg, 60 oblique axial slices of 1 mm in-plane resolution and 3 mm slice thickness). The primary clinical outcome was defined as a composite measure of disability progression which included three clinical assessments: EDSS, Timed 25-Foot Walk (T25FW, a measure of short distance walking speed), and 9-Hole Peg Test (9HPT, a measure of upper limb function) [32]. This composite measure was developed to capture a broader aspect of disability in patients with multiple sclerosis. The 24-week composite confirmed disability progression (CCDP24) was defined as progression on any one of the three components (EDSS, T25FW, or 9HPT): an increase of EDSS score from the baseline at least 1.0 point (or 0.5 points if the baseline EDSS score was larger than 5.5) or a 20% minimum threshold change for T25FW and 9HPT. We also consider a 24-week progression outcome based solely on the 25-foot walking score (S25FW) in Section 3.2.2. Table 1 summarizes the conventional features we use for prediction, grouped by those who experienced CCDP24 progression and those who were censored.
@@
Since MS progression is influenced by many factors, our models consider a variety of demographic and clinical measurements as well as standard imaging metrics. Broadly speaking, we refer to these features as “conventional features” throughout the manuscript to distinguish them from DBM-based features. The conventional features serve as a comparison against the DBM features to assess which feature set performs more optimally. We record nine conventional features which include the patient’s age, birth sex, years of onset, weight (in kg), total brain volume, total T2 lesion volume [^4] The T2-weighted MRI sequence was described by Elliot et al. (2019) [32]: axial 3mm T2- weighted slices were acquired with 2D fast spin-echo, repetition time = 4000–6190 ms, echo time = 74–91 ms and echo train length = 7–11. The T2 lesion volume was measured from T2-weighted MRI images using a proprietary, semi-automatic method. The initial automatic segmentation was performed using a Bayesian classifier [33] which was then verified and corrected by qualified readers., baseline 25-foot walking test score, baseline 9-hole peg test score, and baseline EDSS, which are generally thought to be related to MS disease progression [34].
Table 1: Baseline characteristics stratified by CCDP24 outcome (censored/progressed)
Variable Censored (N = 269) Progressed (N = 81) p-value
Mean (SD) or n / N (%) Mean (SD) or n / N (%)
Baseline age (years) 37 (9) 37 (10) $>0.9$
Years of onset 6.3 (6.2) 7.3 (6.0) 0.088
Baseline weight (kg) 76 (18) 74 (16) 0.4
Baseline 9-hole peg test score time (seconds) 23.1 (6.6) 24.2 (5.2) 0.033
Baseline 25-foot walking test score time (seconds) 7.0 (5.0) 5.9 (3.7) 0.049
Baseline EDSS 2.73 (1.28) 2.57 (1.25) 0.4
T2 lesion volume 8 (9) 13 (14) 0.019
Brain volume 1,501 (86) 1,480 (85) 0.066
Proportion female 185 / 269 (69%) 50 / 81 (62%) 0.2
Wilcoxon rank sum test for continuous variables; Pearson’s Chi-squared test for categorical variable.
@@
Statistical features of regional brain volume were extracted from T1-weighted brain MRI, using a deformation-based morphometry (DBM) pipeline based on the diffeomorphic image registration of advanced normalization tools [35].
@@
The Mindboggle atlas was used to identify individual brain regions [36]. In addition, prior probability images of six major brain regions based on the Mindboggle atlas, including CSF, cortical grey matter, deep grey matter, cortical white matter, the brainstem, and the cerebellum, were binarized at a probability threshold of 0.5. A population-specific group template of the T1-weighted brain MRI was constructed by advanced normalization tools with T1-weighted brain MRI images of 171 healthy adults (aged 20–59 years). Before being fed into the pipeline, each individual brain image was preprocessed with the following steps: 1) resampling to an isotropic resolution of 1 mm; 2) N4 bias correction [37]; and 3) denoising with a non-local algorithm [38]. Regional labels were mapped from the group template to individual images so that volume change of any predefined region could be measured.
As mentioned in the introduction, all of the conventional and DBM features we use for prediction are measured at baseline. The tabular DBM feature set consists of the median and standard deviation voxel values for each of the 51 atlas regions, resulting in 102 total features. Prior to performing feature selection and modeling, we filter out features with a variance below 0.01 across all patients, reducing the number of DBM features from 102 to 56. No other conventional features were excluded in this filtering step.
@@ 2.2 Feature importance using minipatch ensembles
Identifying features that genuinely contribute to prediction is crucial for model interpretability and guiding practical decision-making. Feature selection methods are used to filter out variables that are inconsequential to prediction, retaining those with the most predictive value. Since our dataset consists of high-dimensional and highly correlated features, traditional feature selection methods can potentially misattribute the importance of a feature when other correlated features are present. To offer a more reliable assessment of feature importance, we extend a feature importance method called leave-one-covariate-out minipatch (LOCO-MP) prediction [27, 39, 40, 28] to survival analysis. LOCO-MP consists of taking a subsample of the data, called a “minipatch.” The purple cells in Figure 2 shows a simple example of what three minipatches may look like. A model is then trained on the minipatch and this process is repeated over many minipatches to ensure an even distribution of observations and features are sampled. After training on all minipatches, LOCO-MP determines feature importance for a certain target feature $j$ by comparing the average prediction error across minipatches with and without feature $j$.
LOCO-MP is advantageous over other feature selection techniques due to its ability to provide asymptotic inferential guarantees [28]. Moreover, LOCO-MP can also account for dependencies across features, a common issue in high-dimensional settings like ours – Gan et al. (2023) [28] discuss that, by generating randomly subsampled features across minipatches, LOCO-MP ensures that the predictive value of each feature is not being diminished by other strongly correlated features, since groups of strongly correlated features will not always appear in the same minipatch. The full LOCO-MP procedure and its theoretical guarantees are discussed in Gan et al. (2023) [28].
The procedure is summarized graphically in Figure 2 and outlined below for a target feature $j$:
@@
    Repeat steps 1 and 2 over $K$ minipatches, where $K$ is user-determined.
@@
Our outcome of interest, patient progression, is recorded as a time-to-event pair $(T_{i},C_{i})$, where $T_{i}$ is the event time and $C_{i}\in\left\{0,1\right\}$ is an indicator for whether the patient’s disease progression is observed during the study period ($C_{i}=1$) or is right-censored ($C_{i}=0$). Because our outcomes of interest are recorded in this time-to-event format, we extend LOCO-MP to survival analysis. This is done by predicting discrete instantaneous hazards from the trained models and computing an associated individual-level discrete hazard loss [41, 42]. We refer the reader to A for a full technical description of LOCO-MP and the hazard loss function used to characterize prediction error in our survival models.